find k so that the line joining (3,k) and (2,5) and the joining (-2,3) and (7,0) are parallel.
Answer by mananth(16946)

(3,k) and (2,5) and the joining

(-2,3) and (7,0) are parallel.
slope = 0-3/7-(-2)
slope=-3/9
slope = -1/3.
the slope of the line joining other two points will have same slope since they are parallel.
slope = 5-k / 2-3 = -1/3
5-k / -1 = -1/3
5-k= -1*-1/3
5-k= 1/3
k= 5-1/3
k= 14/3
CHECK
5-14/3 /2-3
=1/3 / -1
=-1/3